OpenStudy (anonymous):

Please help! Medal and fan! The length of one leg of a right triangle is (x - 6) centimeters, and the area is [(1/2)(x^2) - 7x + 24] square centimeters. What is the length of the other leg? I'm pretty sure that the equation i need to solve is [(1/2)(x^2) - 7x + 24) = [(1/2) (x - 6) (h) but I just dont know where to start solving it

OpenStudy (anonymous):

@dtan5457 @Loser66

OpenStudy (anonymous):



OpenStudy (anonymous):

solve for b or h

OpenStudy (wolf1728):

Let's call the other leg "y" so (x-6) * y =[(1/2)(x^2) - 7x + 24] y = [(1/2)(x^2) - 7x + 24] / (x-6)

OpenStudy (anonymous):

1/2*(x^2-7x+24)=1/2bh x^2-7x+24=b(x-6) b=(x2-7x+24)/(x-6)